A positive point charge of magnitude q and mass m is dropped from a great
height h onto another similar charge fixed on the ground. The minimum distance
between the charges in CGS units is:
1
Expert's answer
2019-06-19T10:53:10-0400

 in CGS units the condition of the stop will be

"mg(h - l) = q^{2}(\frac{1}{l} - \frac{1}{h}),"

then

"l = \frac{q^{2}}{mgh}"